# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

### v2

@@ 3.1.2 Document Deshadowing and Document Deblurring
Document deshadowing refers to removing potential shadow within the document images [81]. Document deblurring refers to removing blur artifacts from document images to restore sharp, readable text and clear visual content [82]. Similar to document dewarping, we require the model to output a deshadowed or a deblurred image according to the input image and instruction. The results are presented in Fig. 3. Flux.1-Kontext-dev demonstrates optimal performance in the deshadowing task, which not only removes the shadow but also preserves the original text content and background color. Other models either fail to remove the shadow (OmniGen2, BAGEL), accidentally repeat text content, or change the document’s background color (GPT-4o, Qwen-VLo-Preview). For deblurring, the English case is easier than the Chinese case. Flux.1-Kontext-dev delivers nearly perfect results with clear text and correct content positions, while GPT-4o seems to automatically complete the whole content and generate a new image, possibly witnessing this document’s content during training. OmniGen2 might misunderstand the instruction as erasing all text, thereby delivering a blank image. In the Chinese case, all models can not deblur the text and keep the original content structure (title, subtitle, main text). Surprisingly, although not a unified understanding and generation model, Flux.1-Kontext-dev demonstrates notable low-level visual document processing abilities, showcasing potential for further development.
3.1.3 Document Appearance Enhancement
@@
Document appearance enhancement, also known as document illumination rectification [83, 84], denotes the technique of mainly correcting uneven lighting conditions and removing other real-world degradations like shadows and bleed-through to improve camera-captured documents’ visual quality. Evaluation results are presented in Fig. 4. In the first row, while Flux.1-Kontext-dev achieves relatively good enhancement and text preservation, it adds unexpected woodgrain edges. BAGEL fails to correct the lighting and outputs red-green-mixed background. Qwen-VLo-Preview and Janus-4o produce chaotic, unreadable text. GPT-4o and OmniGen2 exhibit erroneous text repetition similar to Sec.3.1.2. In the second row, no model can perfectly preserve the table content, and the restoration of Chinese text is unsatisfactory. Furthermore, only GPT-4o and Flux.1-Kontext-dev comprehend the instruction to output “PDF-like” documents, suggesting the need to improve generative models’ understanding of professional document terminologies.
3.1.4 Text Editing
We also evaluate the text editing capabilities of existing models on PDF or real-world documents, with results shown in Fig. 5. We discover that GPT-4o, Qwen-VLo-Preview, and Flux.1-Kontext-dev can follow the instruction to perform target modifications, although the modifications are sometimes incomplete or incorrect. However, the last three models consistently fail to edit documents. This may be attributed to their smaller size, i.e., 7B parameters, whereas the first three models possess larger size (The scales of GPT-4o and Qwen-VLo-Preview are unknown but are likely larger than 7B; Flux.1-Kontext-dev has 12B parameters). This indicates that, for text editing on documents, a larger model size is a fundamental requirement. Still, based on a sufficient model scale, further optimization is needed for multiple editing entries and long-sentence editing.
3.2 Historical Document
@@
Figure 6: T2I generation results for historical documents, which typically feature rich and dense text. GPT-4o fulfills generation requirements in most cases with precise English and Chinese text, despite sometimes the content is incomplete or incorrect. However, other models can at most generate ancient book pages, but fail to generate readable textual content. For complete prompts, please refer to our GitHub repository.
@@
Historical documents represent invaluable repositories of human cultural heritage. In recent years, the digitization and analysis of historical documents [85, 86, 87] have emerged as a critical research area in the community, aiming to aid in the preservation and understanding of these ancient cultural artifacts. These documents contain dense textual content, complex layouts, and antiquated fonts, which present great challenge to image generation models. Therefore, we benchmark existing models on various historical document-related tasks, including general tasks like T2I generation and text editing, as well as vertical-domain tasks like historical document restoration.
@@
Figure 7: Text editing and document restoration results for historical documents. For editing, all models fail to modify specific Chinese text. The relatively best GPT-4o suffers from repeated and wrong editing, while the worst Janus-4o can not even output text but unreadable characters. Yet, some models like GPT-4o and Flux.1-Kontext-dev demonstrate promising edit results on English text editing. For historical document restoration, all models fail to generate new text with the original writing style preserved and keep a consistent background.
3.2.1 T2I Generation
@@
We present the T2I generation result of historical documents in Fig. 6. We design long and informative prompts that describe the text content, font style, writing paper details, and other environmental factors like desks and the oil lamp. We ask the model to generate one-page historical scripts in the first to third rows and a three-page script in the fourth row. As observed, only Flux.1-Kontext-dev successfully generates an authentic ancient book in the first row, whereas all other models produce only script-like images reminiscent of ancient writing. This indicates that Flux.1-Kontext-dev may lack historical document training data. However, while models can somehow generate English text, they consistently fail to render Chinese text. GPT-4o is the only model capable of generating both English and Chinese fonts, and also follows the page requirement in the instructions (one-page or three-page). Other models could fail to follow instructions and produce two pages given the one-page requirement. Still, GPT-4o can produce incomplete or incorrect text sometimes.
@@
Figure 8: Results of the style transfer and super resolution tasks for historical documents. For style transfer, while Qwen-VL-Preview produces the relatively best result by preserving the column structure from image 1 and text sparsity from image 2, the rendered text is blurred and lacks semantics. Other models can not follow the instructions and even produce unreadable results. In the super resolution task, GPT-4o and Qwen-VLo-Preview identify the instruction of improving image resolution, while the others can not understand or misunderstand it. Although the resolution is improved, it comes at the cost of color change or font distortion.
3.2.2 Text Editing and Historical Document Restoration
@@
We then perform text editing on historical documents, with results shown in the top frame of Fig. 7. For Chinese text editing, tested models consistently fail to modify Chinese text images. Although GPT-4o locates and modifies the target text, the edited content is incorrect and nearby text is accidentally altered. For English text, GPT-4o and Qwen-VLo-Preview successfully render the target text “COVERING”, but only GPT-4o places it in the correct place. Other models either partially modify the text (Flux.1-Kontext-dev) or fail entirely. Additionally, models introduce unintended changes such as background color alterations and text sharpening effects. This demonstrates that text editing in historical documents remains an unresolved challenge for current models.
@@
Historical document restoration [88] is a technique that aims to recover damaged or deteriorated historical document images, preventing further degradation and restoring their readability. Recently, some methods have explored using diffusion models [89, 90, 91] to accomplish this task. Common approaches require either pre-given missing content [89, 90] or predictions from specialized historical understanding models [91]. For simplicity, we do not provide the missing content and primarily test existing models’ repair ability of text and the damaged area. Results are shown in the bottom frame of Fig. 7. As observed, all models fail to produce text that resembles the original writing style and repair the damaged background texture. They mostly generate a new document rather than conducting “restoration”, demonstrating that existing generative models cannot solve such specialized OCR tasks so far.
3.2.3 Style Transfer and Super Resolution
@@
Additionally, we evaluate the style transfer and super resolution tasks on historical documents, with results presented in Fig. 8. Regarding style transfer, all models fail to meet the target. GPT-4o reversed the target and source styles while generating incorrect content. Surprisingly, despite poor performance in previous tasks, Qwen-VLo-Preview is the only model that preserves the column structure of image 1 and the text sparsity of image 2, achieving the best relative output, although the text remains blurred and unreadable. Other models perform much worse than these two. For super resolution, GPT-4o and Qwen-VLo-Preview recognize the instruction of performing super resolution, but GPT-4o outputs a squared image with changed background color and Qwen-VLo-Preview distorts fonts. Other models seemingly fail to understand or misunderstand the prompt. For example, Flux.1-Kontext-dev actually blurs text rather than clarifying it. Super resolution is a common I2I translation task and we think that current generative models, especially those with trained understanding skills, should be able to perform this basic task. Yet, the results show the opposite and indicate significant room for improvement.
4 Result of Handwritten Text Image
@@
Figure 9: Page-level T2I generation results for handwritten text images. Under the English scenario, most methods can generate accurate text of specific content (although some content may be missed). Qwen-VLo-Preview and Janus-4o are exceptions that generate unreadable and blurred text. However, under the Chinese scenario, only GPT-4o possesses the ability to generate Chinese text, while other models can not recognize the Chinese scripts and produce text in other languages or unreadable text. Flux.1-Kontext-dev even produces unrelated outcomes. For complete prompts, please refer to our GitHub repository.
@@
Handwritten text images are ubiquitous in daily life, appearing in personal notes, letters, and exam papers. Unlike the uniformity of machine-printed fonts, handwriting exhibits significant diversity in style, slant, and hyphenation, reflecting the unique characteristics of individual writers. In recent years, with the rapid advancements in visual generation technologies, the task of handwritten text image generation has garnered substantial attention [63, 92, 93, 94, 95]. This technology not only enables personalized content creation and writing assistance but also provides a new avenue for augmenting training data for handwriting analysis tasks [96, 97, 98, 99, 100, 101]. Furthermore, the editing and removal of text in handwritten images hold broad practical applications in areas such as privacy protection, data cleaning, and educational scenarios [102, 103, 104]. Therefore, in this section, we evaluate the performance and potential of current generative models in three core tasks: handwritten text image generation, text editing, and text removal.
@@
Figure 10: T2I generation results for handwritten text images in other content levels. Similar to the results in Fig. 9, while most models can successfully generate English handwritten text, they generally fail to generate Chinese handwritten lines or characters. Only GPT-4o and Qwen-VLO-Preview demonstrate some successful cases in paragraph or character generation. In addition, all models exhibit the capabilities of handling image-text interleaved scenarios, where GPT-4o performs best. For complete prompts, please refer to our GitHub repository.
4.1 T2I Generation
Generating handwritten text images from textual descriptions is a challenging text-to-image (T2I) generation task. This process demands that a model must accurately render the given textual content while simultaneously adhering to specific stylistic directives, including attributes like handwriting style (e.g., cursive, neat), ink color, and paper type. Consequently, this task serves as a robust benchmark for evaluating a model’s proficiency in both fine-grained text rendering and stylistic control. Our evaluation is structured across multiple levels of granularity: page-level, paragraph-level, line-level, character-level, and interleaved-level.
4.1.1 Page-Level
Fig. 9 presents the results for page-level handwritten text generation. Under the English scenario, most models can generate partially accurate text that is consistent with the prompt. GPT-4o stands out as the relatively best-performing model, producing clear, legible handwriting that closely matches specified styles such as “cursive” or “a slight right slant”. In contrast, Qwen-VLo-Preview can only render a small fraction of correct text, with the majority being unreadable. Janus-4o yields the poorest results, with its output being blurry and entirely illegible.
@@ 4.1.1 Page-Level
Figure 11: Text editing results for handwritten text images. The results show that editing specific handwritten text is a challenging task for existing generation models. They may repeat or lose some text that should be rendered (page-level, paragraph-level) and fail to maintain the original color and background (paragraph-level, line-level). A lot of room is still left for improvement.
The distinction in model capabilities becomes much more pronounced in the Chinese language scenario. Only GPT-4o demonstrates the ability to generate coherent and accurate Chinese text, successfully rendering complex characters in a handwritten style. The other models largely fail to recognize the Chinese script. They either produce unreadable scribbles or default to generating text in other languages. Notably, Flux.1-Kontext-dev exhibits a significant failure by generating a completely unrelated graphic instead of the requested text, indicating a fundamental misunderstanding of the prompt’s intent in a multilingual context.
4.1.2 Other Content Levels
We further evaluate the models on paragraph, line, character, and interleaved text-image generation, with the results presented in Fig. 10. At the paragraph level, most models exhibit significant difficulties with the Chinese paragraph prompt. GPT-4o is the only model to successfully generate a coherent and clear paragraph that accurately renders the text. As the task simplifies to the line and character levels, model performance improves, particularly in the English scenario. For the line generation task, GPT-4o again produces a nearly perfect and clear sentence, followed by OmniGen2. Other models only capture some keywords, while the rest of the text is poorly formatted. At the character level, most models can successfully generate the English letter "P". For the Chinese character, however, only GPT-4o and Qwen-VLo-Preview generate the correct character.
In the interleaved image-text scenario, which requires generating a diagram with explanatory text, GPT-4o demonstrates superior capabilities. It accurately generates a hand-drawn physics diagram of the law of reflection, complete with clear and correctly placed labels. Other models either produce inaccurate diagrams or render the textual labels illegibly.
4.2 Text Editing
We also evaluate the ability of models to edit handwritten text, a task that requires not only accurate understanding of editing instructions but also seamless integration of modifications while preserving the original handwriting style, background, and untouched content.
The performance of different models on text editing tasks is shown in Fig. 11. At the page level, although GPT-4o achieved relatively the best results, it did not perform a true "edit" but instead regenerated the entire page to incorporate the new text, leading to a loss of the original handwriting style and document texture. The performance of other models was even less satisfactory. For example, Qwen-VLo-Preview misinterpreted the prompt and inserted a stamp instead of text; Flux.1-Kontext-dev and OmniGen2 simply output the original text image without adding any new text; BAGEL introduced incorrect text; and Janus-4o produced completely chaotic and unusable results. At the paragraph and line levels, GPT-4o and Qwen-VLo-Preview successfully replaced the text, but failed to maintain the original writing style and untouched content, with issues such as content loss or errors, while other models failed entirely.
In summary, most current models lack the capability for fine-grained, context-aware editing. They often rely on a complete redrawing approach, which cannot retain the necessary style and background attributes of the original image. This highlights significant room for improvement in the task of handwritten text editing.
@@ 4.2 Text Editing
Figure 12: Results of the handwritten text removal task. Comparing the first and third rows, models excel at erasing text in a simple structure image like the case of the first row, while falling short in erasing handwritten text interleaved with other text (e.g., printed text). For removing specific handwriting like in the second and the fourth rows, models usually lack the ability to locate the target text and remove it with other text unchanged.
4.3 Handwritten Text Removal

### v3

@@ 3.2. Document Deblurring
Document deblurring refers to removing blur artifacts from document images to restore sharp, readable text and clear visual content (Souibgui and Kessentini, 2022). Results are presented in the bottom panel of Fig. 3. As observed, English text proves easier to process than Chinese text. Flux.1-Kontext-dev delivers nearly perfect results with clear text and correct content positioning, while GPT-4o seems to automatically complete content and generate a new image, possibly witnessing this document’s content during training. OmniGen2 may misunderstand the instruction as erasing all text, producing blank images. With Chinese text, all models fail to deblur the text while keeping original content structure (title, subtitle, main text). Surprisingly, although not a unified understanding and generation model, Flux.1-Kontext-dev demonstrates notable low-level visual document processing abilities, showcasing potential for further development. In contrast, Qwen-Image outputs drawing-like or unreadable documents, suggesting its shortcomings in processing documents with dense text.
3.2.1. Document Appearance Enhancement
@@
Document appearance enhancement, also known as document illumination rectification (Feng et al., 2021; Zhang et al., 2024a), refers to the technique of mainly correcting uneven lighting conditions while also removing other real-world degradations like shadows and bleed-through to improve camera-captured documents’ visual quality. Evaluation results are presented in Fig. 4. In the first row, while Flux.1-Kontext-dev achieves relatively good enhancement and text preservation, it adds unexpected woodgrain edges. BAGEL fails to correct the lighting and outputs a red-green-mixed background. Qwen-VLo-Preview, Qwen-Image, and Janus-4o generate chaotic, unreadable text. GPT-4o and OmniGen2 exhibit erroneous text repetition similar to the document deshadowing task. In the second row, no model can perfectly preserve the table content, and the restoration of Chinese text is unsatisfactory. Furthermore, only GPT-4o and Flux.1-Kontext-dev comprehend the instruction to output “PDF-like” documents, suggesting the need to improve generative models’ understanding of professional document terminology.
3.2.2. Text Editing
We also evaluate text editing capabilities on PDF or real-world documents, with results shown in Fig. 5. GPT-4o, Qwen-VLo-Preview, and Flux.1-Kontext-dev can follow instructions to perform target modifications, although the modifications are sometimes incomplete or incorrect. However, other models consistently fail to edit documents. OmniGen2 and BAGEL fail to locate the target text and perform no/wrong alterations. Even worse, Qwen-Image and Janus-4o can not preserve the original text content but produce chaotic, unreadable content. As specialized generation models, Flux.1-Kontext-dev significantly outperforms Qwen-Image despite a smaller model scale (13B vs 20B), making it a more feasible option for text-editing fine-tuning in document scenarios due to superior performance and lower training costs.
3.3. Historical Document
@@
Figure 6. T2I generation results for historical documents. GPT-4o fulfills generation requirements of specific page number and the synthesis of both English and Chinese text, despite sometimes the content is incomplete or incorrect. However, other models can generate ancient scripts but mostly fail to produce readable textual content. For complete prompts, please refer to our GitHub repository.
@@
Historical documents represent invaluable repositories of human cultural heritage. The digitization and analysis of historical documents (Philips and Tabrizi, 2020; Nikolaidou et al., 2022; Cao et al., 2024) has emerged as a critical research area, aiming to aid in the preservation and understanding of these ancient cultural artifacts. These documents contain dense textual content, complex layouts, and antiquated fonts, which present a great challenge to image generation models. Therefore, we benchmark existing models on various historical document-related tasks, including general tasks like T2I generation and text editing, as well as vertical-domain tasks like historical document restoration.
@@
Figure 7. Text editing and document restoration results for historical documents. For text editing, all models fail to modify specific Chinese text. The relatively best GPT-4o suffers from repeated and wrong editing, while the worst Qwen-Image even generates new images with specified text content, deviating from the editing objective. Yet, some models like GPT-4o and Flux.1-Kontext-dev demonstrate promising edit results on English text editing. For historical document restoration, all models fail to generate new text while preserving the original writing style and maintaining consistent backgrounds.
3.3.1. T2I Generation
@@
We present T2I generation results for historical documents in Fig. 6. We design detailed prompts describing the text content, font style, writing paper details, and other environmental factors such as desks and oil lamps. We ask models to generate one-page historical scripts in the first three rows and a three-page script in the fourth row. For instruction following, most models successfully generate ancient script, albeit sometimes not one-page or three-page. Conversely, Flux.1-Kontext-dev produces irrelevant outcomes like a book or human figures, suggesting it may lack historical document training data. Regarding text rendering, while models can somehow generate English text, they consistently fail to render Chinese text. GPT-4o is the only model capable of generating both English and Chinese fonts while following page requirements correctly. Nevertheless, GPT-4o sometimes produces incomplete or incorrect text.
@@
Figure 8. Style transfer and super resolution results for historical documents. For style transfer, while Qwen-VL-Preview produces the relatively best result by preserving the column structure from image 1 and text sparsity from image 2, the rendered text is blurred and lacks semantics. Other models can not follow the instructions and produce unreadable results. In the super resolution task, GPT-4o and Qwen-VLo-Preview improve image resolution but at the cost of color change or font distortion, while the others misunderstand the instructions (Flux.1-Kontext-dev) or generate completely unrelated results (Qwen-Image).
3.3.2. Text Editing
@@
We then perform text editing on historical documents, with results shown in the top frame of Fig. 7. For Chinese text editing, tested models consistently fail to modify Chinese text images. Although GPT-4o locates and attempts to modify the target text, the edited content is incorrect and the nearby text is accidentally altered. For English text, GPT-4o and Qwen-VLo-Preview successfully render the target text “COVERING”, but only GPT-4o places it in the correct place. Other models either partially modify the text (Flux.1-Kontext-dev) or fail entirely. Notably, Qwen-Image generates brand-new images rather than editing the original image, indicating its poor editing ability under this scenario. In addition, models introduce unintended changes such as background color alterations and text sharpening effects. This demonstrates that text editing in historical documents remains an unresolved challenge for current models.
3.3.3. Historical Document Restoration
@@
Historical document restoration (Hedjam and Cheriet, 2013) is a technique that aims to recover damaged or deteriorated historical document images, preventing further degradation and restoring readability. Recently, some methods have explored using diffusion models (Zhu et al., 2024b; Yang et al., 2025; Zhang et al., 2025c) for this task. Common approaches require either pre-given missing content (Zhu et al., 2024b; Yang et al., 2025) or predictions from specialized historical understanding models (Zhang et al., 2025c). Here, we do not provide missing content and primarily test existing models’ repair abilities of text and damaged areas. Results are shown in the bottom frame of Fig. 7. All models fail to preserve original writing style or repair damaged background texture. They mostly generate new documents rather than conducting “restoration”, indicating that existing generative models cannot solve such specialized OCR tasks and demand dedicated optimization.
3.3.4. Style Transfer and Super Resolution
@@
Additionally, we evaluate the style transfer and super resolution tasks on historical documents, with results presented in Fig. 8. Regarding style transfer, all models fail to meet the target. GPT-4o reversed the target and source styles while generating incorrect content. Surprisingly, despite poor performance in previous tasks, Qwen-VLo-Preview is the only model that preserves the column structure of image 1 and the text sparsity of image 2, achieving the best relative output, although the text remains blurred and unreadable. Other models perform much worse than these two.
For super resolution, GPT-4o and Qwen-VLo-Preview recognize the instruction of performing super resolution, but GPT-4o outputs a squared image with a changed background color and Qwen-VLo-Preview distorts fonts. Other models misunderstand or entirely fail to comprehend the prompt. Flux.1-Kontext-dev actually blurs text instead of enhancing clarity, while Qwen-Image generates an unrelated map image rather than improving resolution, further demonstrating its limitations in image editing tasks. Super resolution is a common image-to-image translation task and we think that current generative models, especially those with trained understanding skills, should be able to perform this basic task. Yet, the results reveal significant deficiencies and highlight substantial room for improvement in existing models.
4. Result of Handwritten Text Image
@@
Figure 9. Page-level T2I generation results for handwritten text images. For English, most methods can generate accurate texts of specific content (although some content may be missed). Qwen-VLo-Preview and Janus-4o are exceptions that generate unreadable and blurred text. However, for Chinese, GPT-4o and Qwen-Image exhibit the ability to generate Chinese text, while other models can not recognize Chinese scripts and produce text in other languages or unreadable text. Flux.1-Kontext-dev even produces unrelated outcomes. For complete prompts, please refer to our GitHub repository.
@@
Handwritten text images are ubiquitous in personal notes, letters, and exam papers. Unlike the uniformity of machine-printed fonts, handwriting exhibits significant diversity in style, slant, and hyphenation, reflecting the unique characteristics of individual writers. Recent advancements in visual generation technologies have brought increasing attention to handwritten text image generation (Zhu et al., 2023; Das et al., 2023; Dai et al., 2024; Pippi et al., 2025). This technology enables writing assistance (Ha and Eck, 2018; Schäfer et al., 2023) and augments training data for handwriting analysis tasks (Luo et al., 2019; Zhang et al., 2022a; Peng et al., 2022; Yang et al., 2023; Zhang and Jin, 2024; Zhang et al., 2025a). Additionally, the editing and removal of text in handwritten images hold broad practical applications in areas such as privacy protection, data cleaning, and educational scenarios (Li et al., 2024b; Huang et al., 2023; Wang et al., 2022). We therefore evaluate the performance of current models in three tasks: handwritten text image generation, text editing, and text removal.
@@
Figure 10. T2I generation results for handwritten text images in other content levels. Similar to Fig. 9, while most models successfully generate English handwritten text, they generally fail with Chinese handwritten lines or characters. Only GPT-4o and Qwen-VLo-Preview demonstrate some successful cases in paragraph or character generation. In addition, all models exhibit capabilities for handling image-text interleaved scenarios, with GPT-4o performing best. For complete prompts, please refer to our GitHub repository.
@@
Figure 11. Text editing results for handwritten text images. The results show that editing specific handwritten text is a challenging task for existing generation models. They may repeat or lose some text that should be rendered (page-level, paragraph-level) and fail to maintain the original color and background (paragraph-level, line-level). A lot of room is still left for improvement.
@@
Figure 12. Text removal results for handwritten text images. Comparing the first and third rows, models excel at erasing text in simple structure images, while falling short in erasing handwritten text interleaved with other text (e.g., printed text). For removing specific handwriting, as shown in the second and fourth rows, models typically fail to accurately locate target text while preserving surrounding content.
4.1. T2I Generation
Generating handwriting images with textual descriptions is a challenging T2I generation task. It demands models to accurately render given textual content while adhering to specific stylistic directives, such as writing style, ink color, and paper type. We evaluate this task across multiple granularity levels: page-level, paragraph-level, line-level, character-level, and interleaved-level.
4.1.1. Page-Level
Fig. 9 presents the results for page-level handwritten text generation. Under the English scenario, most models can generate partially accurate text consistent with the prompt. GPT-4o performs best, producing clear, legible handwriting that closely matches specified styles such as “cursive” or a “slight right slant”. Qwen-Image also performs well by producing accurate text, albeit it does not follow the specified writing style instructions. In contrast, Qwen-VLo-Preview can only render a small fraction of correct text, with the majority being unreadable. Janus-4o yields the poorest results, with its output being blurry and entirely illegible.
The distinction in model capabilities becomes much more pronounced in the Chinese language scenario. GPT-4o and Qwen-Image demonstrate the ability to generate coherent and accurate Chinese handwritten text. The other models largely fail to recognize the Chinese script. They either produce totally unreadable scribbles or default to generating text in other languages. Notably, Flux.1-Kontext-dev exhibits a significant failure by generating a completely irrelevant graphic instead of the requested text, indicating the inability to perceive the prompt’s intent.
4.1.2. Other Content Levels
We further evaluate the models on paragraph, line, character, and interleaved handwritten text generation, with results presented in Fig. 10. At the paragraph level, most models struggle with Chinese prompts. Only GPT-4o successfully generates coherent, clear paragraphs with accurately rendered text. Model performance improves as tasks simplify to line and character levels, particularly in the English scenario. For line generation, GPT-4o and Qwen-Image produce correct and clear sentences, followed by OmniGen2 with nearly perfect output. Other models only capture some keywords, while the rest of the text is poorly formatted. At the character level, most models can successfully generate the English letter “P”. For Chinese characters, however, only GPT-4o, Qwen-VLo-Preview, and Qwen-Image generate correct contents. For interleaved image-text generation, which requires generating a diagram with explanatory text, GPT-4o demonstrates superior capabilities. It accurately generates a hand-drawn physics diagram of the law of reflection, complete with clear and correctly placed labels. Other models either produce inaccurate diagrams or render the textual labels illegible.
4.2. Text Editing
We perform text editing on handwritten text images, with results shown in Fig. 11. At the page level, although GPT-4o achieved relatively optimal results, the output image is inadvertently cropped to a square shape and some content is lost. Other models perform even less satisfactorily. For example, Qwen-VLo-Preview misinterprets the prompt and inserts a stamp with incomplete text. Qwen-Image replaces the title with the target text but does not produce the embossed effect. Flux.1-Kontext-dev and OmniGen2 simply output the original image without any alteration. At paragraph and line levels, GPT-4o and Qwen-VLo-Preview successfully replace text but fail to maintain the original writing style and content, with issues such as content loss or errors, while other models fail entirely. The results reveal that most current models lack the capability for fine-grained editing. They often rely on a complete redrawing approach, which cannot retain the necessary style and background attributes of the original image. This highlights significant room for improvement in this task.
4.3. Handwritten Text Removal